OpenStudy (anonymous):

solve the equation. Give a general formula for all the solutions. 2 cosθ + 1 = 0

OpenStudy (anonymous):

\[2\cos \theta+1=0\] \[2\cos \theta= -1\] \[\cos \theta=-\frac{ 1 }{ 2 }\] \[\cos \theta=-\cos \frac{ \pi }{ 3 }\] \[\cos \theta=\cos(\pi+\frac{ \pi }{ 3 })\] \[\theta=2n \pi \pm \frac{ \pi }{ 3 }\] ...where \[n \epsilon I\]

OpenStudy (anonymous):

The answer choice are A.) θ=3π/2+kπ B.) θ=2π/3+2kπ, θ=4π/3+2kπ C.) θ=π/2+2kπ, θ=3π/2+2kπ D.) θ=2π/3+kπ, θ=4π/3+kπ

OpenStudy (anonymous):

hey sry I was wrong in last step \[\theta=2n \pi+\frac{ 4\pi }{ 3 }\]..........(1) and \[\theta=2n \pi-\frac{ 4\pi }{ 3 }\] ........(2) where nϵI now put n=k in equation 1 and n= k+1 in equation 2 you will get the answer

OpenStudy (anonymous):

That's fine. Thank you so much ^^